the speed of a boat in still water is 15 km/hr. it needs four more hours to travel 63 km against the current of a river than it needs to travel down the river. Determine the speed of the current of the river.
Answer by scott8148(6628)

[63 / (15 - x)] + 4 = 63 / (15 + x)

multiplying by LCD (15^2 - x^2) ___ [63 (15 + x)] + [4 (15^2 - x^2)] = 63 (15 - x)